Papers by Andrew C. Haas

1. “Limits on anomalous trilinear gauge couplings from W W → e^{+}e^{−}, W W → e^{±}μ^{∓}, and W W → μ^{+}μ^{−} events from pp̄ collisions at \sqrt{s} = 1.96-TeV” V. M. Abazov et al. [D0 Collaboration] Phys. Rev. D 74, 057101 (2006) [arXiv:hep-ex/0608011]

2. “Search for scalar leptoquarks in the acoplanar jet topology in p anti-p collisions at s**(1/2) = 1.96-TeV” V. M. Abazov et al. [D0 Collaboration] Phys. Lett. B 640, 230 (2006) [arXiv:hep-ex/0607009]

3. “Search for a heavy resonance decaying into a Z + jet final state in p anti-p collisions at s**(1/2) = 1.96-TeV using the D0 detector” V. M. Abazov et al. [D0 Collaboration] Phys. Rev. D 74, 011104 (2006) [arXiv:hep-ex/0606018]

4. “Search for resonant second generation slepton production at the Tevatron”

V. M. Abazov et al. [D0 Collaboration] Phys. Rev. Lett. 97, 111801 (2006) [arXiv:hep-ex/0605010]

5. “Search for neutral Higgs bosons decaying to tau pairs in p anti-p collisions at s**(1/2) = 1.96-TeV” V. M. Abazov et al. [D0 Collaboration] Phys. Rev. Lett. 97, 121802 (2006) [arXiv:hep-ex/0605009]

6. “Search for R-parity violating supersymmetry via the LLE couplings lambda(121), lambda(122) or lambda(133) in p anti-p collisions at s**(1/2) = 1.96-TeV” V. M. Abazov et al. [D0 Collaboration] Phys. Lett. B 638, 441 (2006) [arXiv:hep-ex/0605005]

7. “A precise measurement of the B/s0 lifetime”

V. M. Abazov et al. [D0 Collaboration] arXiv:hep-ex/0604046 (Submitted to Phys.Rev.Lett.)

8. “Search for excited muons in p anti-p collisions at s**(1/2) = 1.96-TeV”

V. M. Abazov et al. [D0 Collaboration] Phys. Rev. D 73, 111102 (2006) [arXiv:hep-ex/0604040]

9. “Search for squarks and gluinos in events with jets and missing transverse energy in p anti-p collisions at s**(1/2) = 1.96-TeV” V. M. Abazov et al. [D0 Collaboration] Phys. Lett. B 638, 119 (2006) [arXiv:hep-ex/0604029]

10. “Multivariate searches for single top quark production with the D0 detector” V. M. Abazov et al. [D0 Collaboration] arXiv:hep-ex/0604020 (Submitted to Phys.Rev.D)

11. “Search for the rare decay B0/s –¿ Phi mu+ mu- with the D0 detector” V. M. Abazov et al. [D0 Collaboration] Phys. Rev. D 74, 031107 (2006) [arXiv:hep-ex/0604015]

12. “First direct two-sided bound on the B/s0 oscillation frequency”

V. M. Abazov et al. [D0 Collaboration] Phys. Rev. Lett. 97, 021802 (2006) [arXiv:hep-ex/0603029] 13. “Measurement of B(t –¿ W b)/B(t –¿ W q) at s**(1/2) = 1.96-TeV”

V. M. Abazov et al. [D0 Collaboration] Phys. Lett. B 639, 616 (2006) [arXiv:hep-ex/0603002]

14. “Search for pair production of second generation scalar leptoquarks in p anti-p collisions at s**(1/2) = 1.96-TeV” V. M. Abazov et al. [D0 Collaboration] Phys. Lett. B 636, 183 (2006) [arXiv:hep-ex/0601047]

15. “Measurement of the isolated photon cross section in p anti-p collisions at s**(1/2) = 1.96-TeV” V. M. Abazov et al. [D0 Collaboration] Phys. Lett. B 639, 151 (2006) [arXiv:hep-ex/0511054]

16. “Search for the Higgs boson in H –¿ W W(*) decays in p anti-p collisions at s**(1/2) = 1.96-TeV” V. M. Abazov et al. [D0 Collaboration] Phys. Rev. Lett. 96, 011801 (2006) [arXiv:hep-ex/0508054]

17. “The upgraded D0 detector”

V. M. Abazov et al. [D0 Collaboration] Nucl. Instrum. Meth. A 565, 463 (2006) [arXiv:physics/0507191]

18. “Measurement of the lifetime difference in the B/s0 system”

V. M. Abazov et al. [D0 Collaboration] Phys. Rev. Lett. 95, 171801 (2005) [arXiv:hep-ex/0507084]

19. “Measurement of semileptonic branching fractions of B mesons to narrow D** states” V. M. Abazov et al. [D0 Collaboration] Phys. Rev. Lett. 95, 171803 (2005) [arXiv:hep-ex/0507046]

20. “Search for large extra spatial dimensions in dimuon production at D0”

V. M. Abazov et al. [D0 Collaboration] Phys. Rev. Lett. 95, 161602 (2005) [arXiv:hep-ex/0506063]

21. “Measurement of the t anti-t production cross section in p anti-p collisions at s**(1/2) = 1.96-TeV in dilepton final states” V. M. Abazov et al. [D0 Collaboration] Phys. Lett. B 626, 55 (2005) [arXiv:hep-ex/0505082]

22. “Search for single top quark production in p anti-p collisions at s**(1/2) = 1.96-TeV” V. M. Abazov et al. [D0 Collaboration] Phys. Lett. B 622, 265 (2005) [arXiv:hep-ex/0505063]

23. “Measurement of the W boson helicity in top quark decays”

V. M. Abazov et al. [D0 Collaboration] Phys. Rev. D 72, 011104 (2005) [arXiv:hep-ex/0505031]

24. “Search for Randall-Sundrum gravitons in dilepton and diphoton final states” V. M. Abazov et al. [D0 Collaboration] Phys. Rev. Lett. 95, 091801 (2005) [arXiv:hep-ex/0505018]

25. “Measurement of the t anti-t production cross section in p anti-p collisions at s**(1/2) = 1.96-TeV using lepton + jets events with lifetime b-tagging” V. M. Abazov et al. [D0 Collaboration] Phys. Lett. B 626, 35 (2005) [arXiv:hep-ex/0504058]

26. “Measurement of the t anti-t production cross section in p anti-p collisions at s**(1/2) = 1.96-TeV using kinematic characteristics of lepton + jets events” V. M. Abazov et al. [D0 Collaboration] Phys. Lett. B 626, 45 (2005) [arXiv:hep-ex/0504043] 27. “Search for supersymmetry via associated production of charginos and neutralinos in

final states with three leptons” V. M. Abazov et al. [D0 Collaboration] Phys. Rev. Lett. 95, 151805 (2005) [arXiv:hep-ex/0504032]

28. “Production of W Z events in p anti-p collisions at s**(1/2) = 1.96-TeV and limits on anomalous W W Z couplings” V. M. Abazov et al. [D0 Collaboration] Phys. Rev. Lett. 95, 141802 (2005) [arXiv:hep-ex/0504019]

29. “Search for neutral supersymmetric Higgs bosons in multijet events at s**(1/2) = 1.96-TeV” V. M. Abazov et al. [D0 Collaboration] Phys. Rev. Lett. 95, 151801 (2005) [arXiv:hep-ex/0504018]

30. “Measurement of the p anti-p –¿ W gamma + X cross section at s**(1/2) = 1.96-TeV and W W gamma anomalous coupling limits” V. M. Abazov et al. [D0 Collaboration] Phys. Rev. D 71, 091108 (2005) [arXiv:hep-ex/0503048]

31. “Study of Z gamma events and limits on anomalous Z Z gamma and Z gamma gamma couplings in p anti-p collisions at s**(1/2) = 1.96-TeV” V. M. Abazov et al. [D0 Collaboration] Phys. Rev. Lett. 95, 051802 (2005) [arXiv:hep-ex/0502036]

32. “Measurement of inclusive differential cross sections for Upsilon(1S) production in p antip collisions at s**(1/2) = 1.96-TeV” V. M. Abazov et al. [D0 Collaboration] Phys. Rev. Lett. 94, 232001 (2005) [arXiv:hep-ex/0502030]

33. “The DZERO level 3 data acquisition system”

R. D. Angstadt et al. IEEE Trans. Nucl. Sci. 51, 445 (2004)

34. “Search for first-generation scalar leptoquarks in p anti-p collisions at s**(1/2) = 1.96-TeV” V. M. Abazov et al. [D0 Collaboration] Phys. Rev. D 71, 071104 (2005) [arXiv:hep-ex/0412029]

35. “First measurement of sigma(p anti-p –¿ Z) x Br(Z –¿ tau tau) at s**(1/2) = 1.96-TeV” V. M. Abazov et al. [D0 Collaboration] Phys. Rev. D 71, 072004 (2005) [arXiv:hep-ex/0412020]

36. “A search for anomalous heavy-flavor quark production in association with W bosons” V. M. Abazov et al. [D0 Collaboration] Phys. Rev. Lett. 94, 152002 (2005) [arXiv:hep-ex/0411084]

37. “A measurement of the ratio of inclusive cross sections σ(pp → Z + b − jet)/σ(pp → Z + jet) at \sqrt{s} = 1.96 TeV” V. M. Abazov et al. [D0 Collaboration] Phys. Rev. Lett. 94, 161801 (2005) [arXiv:hep-ex/0410078]

38. “Measurement of the W W production cross section in pp̄ collisions at \sqrt{s} = 1.96 TeV” V. M. Abazov et al. [D0 Collaboration] Phys. Rev. Lett. 94, 151801 (2005) [arXiv:hep-ex/0410066]

39. “A search for W bb̄ and W H production in pp̄ collisions at \sqrt{s} = 1.96 TeV”

V. M. Abazov et al. [D0 Collaboration] Phys. Rev. Lett. 94, 091802 (2005) [arXiv:hep-ex/0410062]

40. “Measurement of the Λ_{b}^{0} lifetime in the decay Λ_{b}^{0} → J/ψΛ^{0} with the DØ detector” V. M. Abazov et al. [D0 Collaboration] Phys. Rev. Lett. 94, 102001 (2005) [arXiv:hep-ex/0410054] 41. “Measurement of the ratio of B^{+} and B^{0} meson lifetimes”

V. M. Abazov et al. [D0 Collaboration] Phys. Rev. Lett. 94, 182001 (2005) [arXiv:hep-ex/0410052]

42. “A search for the flavor-changing neutral current decay B_{s}^{0} → μ^{+}μ^{−} in pp̄ collisions at \sqrt{s} = 1.96 TeV with the DØ detector” V. M. Abazov et al. [D0 Collaboration] Phys. Rev. Lett. 94, 071802 (2005) [arXiv:hep-ex/0410039]

43. “Measurement of the B_{s}^{0} lifetime in the exclusive decay channel B_{s}^{0} → J/ψϕ” V. M. Abazov et al. [D0 Collaboration] Phys. Rev. Lett. 94, 042001 (2005) [arXiv:hep-ex/0409043]

44. “Measurement of dijet azimuthal decorrelations at central rapidities in pp̄ collisions at \sqrt{s} = 1.96 TeV” V. M. Abazov et al. [D0 Collaboration] Phys. Rev. Lett. 94, 221801 (2005) [arXiv:hep-ex/0409040]

45. “Search for supersymmetry with gauge-mediated breaking in diphoton events at DØ” V. M. Abazov et al. [D0 Collaboration] Phys. Rev. Lett. 94, 041801 (2005) [arXiv:hep-ex/0408146]

46. “Observation and properties of the X(3872) decaying to J/ψπ^{+}π^{−} in pp̄ collisions at \sqrt{s} = 1.96 TeV” V. M. Abazov et al. [D0 Collaboration] Phys. Rev. Lett. 93, 162002 (2004) [arXiv:hep-ex/0405004]

47. “Search for doubly-charged Higgs boson pair production in the decay to μ^{+}μ^{+}μ^{−}μ^{−} in pp̄ collisions at \sqrt{s} = 1.96 TeV” V. M. Abazov et al. [D0 Collaboration] Phys. Rev. Lett. 93, 141801 (2004) [arXiv:hep-ex/0404015]

48. “A search for neutral Higgs bosons at high tan beta in multi-jet events from p anti-p collisions at s**(1/2) = 1960-GeV” A. C. Haas FERMILAB-THESIS-2004-26

49. “The DZERO level 3 data acquisition system”

D. Chapin et al. eConf C0303241, TUGP010 (2003) [eConf C0303241, MOGT002 (2003)] [arXiv:physics/0307070]

50. “D0 online monitoring and automatic DAQ recovery”

A. Haas et al. eConf C0303241, TUGP011 (2003) [eConf C0303241, THGT004 (2003)] [arXiv:physics/0306195]

51. “Simultaneous tracking and vertexing with elastic templates”

A. Haas Prepared for 7th International Workshop on Advanced Computing and Analysis Techniques in Physics Research (ACAT 2000), Batavia, Illinois, 16-20 Oct 2000